# Entanglement generation and scaling from noisy quenches across a quantum critical point

### v2

@@ V.2 Computing concurrence for time-evolved states of the TFI chain
where, in the fourth equation, we have dropped the subindex $j$ and used $\mbox{Tr}\otimes_{k}\rho_{k}=\prod_{k}\mbox{Tr}\,\rho_{k}=1$. The replacement of $(p,q)$ by $(k_{j},-k_{j})$ in the third equation follows from the fact that each Fock space, labelled by $k_{j}$, is spanned by states with 0 and 2 quasiparticles, implying that $p=-q=k_{j}$ (with $k_{j}$ summed over); cf. Eq. (II).
To take the trace in the last line of Eq. (V.2) we use the diagonal bases $\{|\phi^{\pm}_{k}\}_{k}$ of the mode Hamiltonians at fixed time $t$,
$\mbox{Tr}(\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger})=\langle\phi^{-}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{-}_{k}\rangle+\langle\phi^{+}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{+}_{k}\rangle$. (A32)
@@
$\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle=\frac{1}{N}\sum_{k>0}\Big{[}\sin(kr)\big{(}\sin(2\theta_{k})(\rho^{(d)}_{k,22}-\rho^{(d)}_{k,11})+2i(\cos^{2}(\theta_{k})\,\rho^{(d)}_{k,12}+\sin^{2}(\theta_{k})\,\rho^{(d)}_{k,21})\big{)}\Big{]}$. (A34)
@@
$\langle c_{\l }c_{\l +r}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big{[}\sin(kr)\big{(}\sin(2\theta_{k})(\rho^{(d)}_{k,11}-\rho^{(d)}_{k,22})+2i(\cos^{2}(\theta_{k})\,\rho^{(d)}_{k,21}+\sin^{2}(\theta_{k})\,\rho^{(d)}_{k,12})\big{)}\Big{]}$, (A35)
$\langle c_{\l }^{\dagger}c_{\l +r}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big{[}\cos(kr)\big{(}2\cos^{2}(\theta_{k})\rho^{(d)}_{k,22}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,11}-i\sin(2\theta_{k})(\rho^{(d)}_{k,12}-\rho^{(d)}_{k,21})\big{)}\Big{]}$, (A36)
$\langle c_{\l }c_{\l +r}^{\dagger}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big{[}\cos(kr)\big{(}2\cos^{2}(\theta_{k})\rho^{(d)}_{k,11}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,22}-i\sin(2\theta_{k})(\rho^{(d)}_{k,21}-\rho^{(d)}_{k,12})\big{)}\Big{]}$. (A37)
By feeding the expressions from Eqs. (A34)-(A37) into the formulas in Eq. (V.2), we finally obtain the $A$ and $B$ two-point functions that determine the spin correlations in Eqs. (V.2) and (V.2). Pruning the outcome, the result is
@@
$\langle B_{l}B_{l+r}\rangle$ $=$ $\frac{2i}{N}\sum_{k>0}\mbox{Re}(\rho^{(d)}_{k,12})\sin(kr)-\delta_{r,0},$ (A38)
@@
In the thermodynamic limit $N\rightarrow\infty$ this can be written as
@@
Also note that for a quench from $h(t_{\text{initial}})\!\rightarrow\!$-$\infty$ to $h(t_{\text{final}})\!\rightarrow\!+\infty$, the off-diagonal terms $\rho^{(d)}_{k,12}(t)=(\rho^{(d)}_{k,21})^{\ast}$ of the density matrix $\rho_{k}(t)$ vanish, as does $\theta_{k}(t)$. As a consequence, the two-point fermion functions in Eq. (V.2) reduce to
@@
$\langle A_{l}B_{l+r}\rangle$ $=$ $-\langle B_{l+r}A_{l}\rangle=\frac{1}{N}\sum_{k}(1-2\rho^{(d)}_{k,22})\cos(kr),$ (A39)
@@
$\langle A_{l}B_{l+r}\rangle$ $=$ $-\langle B_{l+r}A_{l}\rangle=\frac{1}{2\pi}\int_{\pi}^{\pi}(1-2p_{k}(t))\cos(kr)dk=-\frac{2}{\pi}\int_{0}^{\pi}p_{k}(t)\cos(kr)dk,$ (A40)
@@
Returning to our task to compute the concurrences $C_{l,l+1}(t)$ and $C_{l,l+2}(t)$ for nearest- and next-nearest-neighbor spins generated by a ramped quench, we now have all the pieces in place. From Eq. (V.2) we obtain closed expressions for the spin correlations in Eqs. (V.2) and (V.2), as well as for the one-point function in Eq. (A11). Eq. (V.2) then yields the elements of the reduced density matrices $\rho_{l,l+1}$ and $\rho_{l,l+2}$, cf. Eq. (A7). With this, expressions for the concurrences are obtained from Eqs. (A1) and (A2). What remains is to compute the time-dependent matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$ and $\rho^{(d)}_{k,22}(t)$ that appear in Eq. (V.2) (with the suppressed time argument reinstated). For the noiseless pure ensemble, this can be done by numerically integrating the von Neumann equation in the diagonal basis,
$\dot{\rho}_{0,k}(t)=-i[H_{0,k}(t),\rho_{0,k}(t)],$ (A41)
for all $k$, with $t$ the time when the quench is completed. As initial condition we impose that the system is in the noiseless ground state at time $t_{i}$, i.e., that all modes $k$ occupy the lower levels $|\phi_{k}^{-}(t_{i})\rangle$. As expected for a pure ensemble, note that the matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$ and $\rho^{(d)}_{k,22}(t)$ defined in Eq. (V.2) (with the time argument suppressed) map onto those in Eq. (III), where in the latter equation the density operator for mode $k$ is denoted $\rho_{0,k}(t)$ while in Eq. (V.2) the more general symbol $\rho_{k}$ - adaptable to pure as well as mixed ensembles - is used.
@@
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}[\sigma^{z},[\sigma^{z},\rho_{k}(t)]],$ (A42)
where $\rho_{k}(t)=\langle\rho_{\eta,k}(t)\rangle$ is the averaged density matrix for mode $k$ in the presence of white noise, with the average taken over the white-noise distribution $\{\eta(t)\}$ specified in Eq. (10). By numerically integrating the master equation, Eq. (A42), for all $k$, with the same initial condition as in the noiseless case, we extract the matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$ and $\rho^{(d)}_{k,22}(t)$ in the diagonal basis. In this way we finally obtain the mean concurrences $\langle C_{l,l+1}(t)\rangle$ and $\langle C_{l,l+2}(t)\rangle$ for nearest- and next-nearest-neighbor spins.
@@
V.3 Noise-averaged nonequilibrium concurrence: exact master equation
@@
Let $\rho(t)=\langle\rho_{\eta}(t)\rangle$ be the ensemble average over many noise realizations (all with a common noise-free initial condition). The averaged von Neumann equation (A44) then takes the form
$\dot{\rho}(t)=-i[H_{0}(t),\rho(t)]-i[H_{1}(t),\langle\eta(t)\rho_{\eta}(t)\rangle].$ (A46)
@@
$\langle\eta(t)\rho_{\eta}(t)\rangle=\langle\eta(t)\rangle\langle\rho_{\eta}(t)\rangle+\int_{0}^{t}\langle\eta(t)\eta(s)\rangle\langle\frac{\delta\rho_{\eta}}{\delta\eta}\rangle ds,$ (A47)
@@
$\frac{\delta\rho_{\eta}}{\delta\eta}=\frac{\partial\dot{\rho}_{\eta}}{\partial\eta}-\frac{\mbox{d}}{\mbox{d}t}\frac{\partial\dot{\rho}_{\eta}}{\partial\dot{\eta}}.$ (A48)
Combining Eqs. (A48), (A43), and (A44) gives
$\frac{\delta\rho_{\eta}}{\delta\eta}=-i[H_{1}(t),\eta(t)].$ (A49)
The master equation follows by inserting Eq. (A47) with $\langle\eta(t)\rangle=0$ into (A46), using Eqs. (10) and (A49),
$\dot{\rho}(t)=-i[H_{0}(t),\rho(t)]-\frac{\xi^{2}}{2}\Big{[}H_{1}(t),[H_{1}(t),\rho(t)]\Big{]},$ (A50)
where the factor of $1/2$ multiplying the noise term comes from the delta function in Eq. (10) having support at the edge of the integral in Eq. (A47).
As follows from the discussion in Sec. II, the TFI Hamiltonian $H_{0}(t)$ in Eq. (1) with an added white-noise term gets expressed as a sum over decoupled mode Hamiltonians $H_{k}(t)=H_{0,k}(t)+\eta(t)H_{1}$ after a Jordan-Wigner transformation, with $H_{0,k}(t)$ given in Eq. (4) and with $H_{1}=\sigma^{z}$. As a consequence, the density matrix $\rho_{\eta}(t)$ corresponding to $H_{0}(t)$ with the added white-noise term has a direct product structure, i.e., $\rho_{\eta}(t)=\otimes_{k}\rho_{k,\eta}(t)$ with the $2\times 2$ density matrix $\rho_{k,\eta}(t)$ satisfying $\dot{\rho}_{k,\eta}(t)=-i[H_{0,k}(t),\rho_{k,\eta}(t)]$ for a single realization of the noise function $\eta$, common to all modes $k$. The master equation for the noise-averaged density matrix $\rho_{k}(t)=\langle\rho_{k,\eta}(t)\rangle$ is read off from Eq. (A50), using that $H_{1}=\sigma^{z}$, with the result
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}\Big{[}\sigma^{z},[\sigma^{z},\rho_{k}(t)]\Big{]}.$ (A51)

### v3

@@ Acknowledgement
where, in the fourth equation, we have dropped the subindex $j$ and used $\mbox{Tr}\otimes_{k}\rho_{k}=\prod_{k}\mbox{Tr}\,\rho_{k}=1$. The replacement of $(p,q)$ by $(k_{j},-k_{j})$ in the third equation follows from the fact that each Fock space, labeled by $k_{j}$, is spanned by two states with zero and two quasiparticles, respectively, implying that $p=-q=k_{j}$ (with $k_{j}$ summed over) [cf. Eq. (II)].
To take the trace in the last line of Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) we use the diagonal bases $\{|\phi^{\pm}_{k}\}_{k}$ of the mode Hamiltonians at fixed time $t$,
$\mbox{Tr}(\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger})=\langle\phi^{-}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{-}_{k}\rangle+\langle\phi^{+}_{k}|\rho_{k}c_{k}^{\dagger}c_{-k}^{\dagger}|\phi^{+}_{k}\rangle$. (B32)
@@
$\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle=\frac{1}{N}\sum_{k>0}\Big[\sin(kr)\big(\sin(2\theta_{k})(\rho^{(d)}_{k,22}-\rho^{(d)}_{k,11})+2i(\cos^{2}(\theta_{k})\,\rho^{(d)}_{k,12}+\sin^{2}(\theta_{k})\,\rho^{(d)}_{k,21})\big)\Big$]. (B34)
@@
$\langle c_{\l }c_{\l +r}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big[\sin(kr)\big(\sin(2\theta_{k})(\rho^{(d)}_{k,11}-\rho^{(d)}_{k,22})+2i(\cos^{2}(\theta_{k})\,\rho^{(d)}_{k,21}+\sin^{2}(\theta_{k})\,\rho^{(d)}_{k,12})\big)\Big$], (B35)
$\langle c_{\l }^{\dagger}c_{\l +r}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big[\cos(kr)\big(2\cos^{2}(\theta_{k})\rho^{(d)}_{k,22}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,11}-i\sin(2\theta_{k})(\rho^{(d)}_{k,12}-\rho^{(d)}_{k,21})\big)\Big$], (B36)
$\langle c_{\l }c_{\l +r}^{\dagger}\rangle$ $=$ $\frac{1}{N}\sum_{k>0}\Big[\cos(kr)\big(2\cos^{2}(\theta_{k})\rho^{(d)}_{k,11}+2\sin^{2}(\theta_{k})\rho^{(d)}_{k,22}-i\sin(2\theta_{k})(\rho^{(d)}_{k,21}-\rho^{(d)}_{k,12})\big)\Big$]. (B37)
By feeding the expressions from Eqs. (B34)-(B37) into the formulas in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain), we finally obtain the $A$ and $B$ two-point functions that determine the spin correlations in Eqs. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) and (Appendix B: Computing concurrence for time-evolved states of the TFI chain). Pruning the outcome, the result is
@@
$\langle B_{l}B_{l+r}\rangle$ $=$ $\frac{2i}{N}\sum_{k>0}\mbox{Re}(\rho^{(d)}_{k,12})\sin(kr)-\delta_{r,0},$ (B38)
@@
In the thermodynamic limit $N\rightarrow\infty$, this can be written as
@@
Also note that for a quench from $h(t_{\text{initial}})\!\rightarrow\!$-$\infty$ to $h(t_{\text{final}})\!\rightarrow\!+\infty$, the off-diagonal terms $\rho^{(d)}_{k,12}(t)=(\rho^{(d)}_{k,21})^{\ast}$ of the density matrix $\rho_{k}(t)$ vanish, as does $\theta_{k}(t)$. As a consequence, the two-point fermion functions in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) reduce to
@@
$\langle A_{l}B_{l+r}\rangle$ $=$ $-\langle B_{l+r}A_{l}\rangle=\frac{1}{N}\sum_{k}(1-2\rho^{(d)}_{k,22})\cos(kr),$ (B39)
@@
$\langle A_{l}B_{l+r}\rangle$ $=$ $-\langle B_{l+r}A_{l}\rangle=\frac{1}{2\pi}\int_{\pi}^{\pi}(1-2p_{k}(t))\cos(kr)dk=-\frac{2}{\pi}\int_{0}^{\pi}p_{k}(t)\cos(kr)dk,$ (B40)
@@
Returning to our task to compute the concurrences $C_{l,l+1}(t)$ and $C_{l,l+2}(t)$ for nearest- and next-nearest-neighbor spins generated by a ramped quench, we now have all the pieces in place. From Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain), we obtain closed expressions for the spin correlations in Eqs. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) and (Appendix B: Computing concurrence for time-evolved states of the TFI chain), as well as for the one-point function in Eq. (B11). Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) then yields the elements of the reduced density matrices $\rho_{l,l+1}$ and $\rho_{l,l+2}$ [cf. Eq. (B7)]. With this, expressions for the concurrences are obtained from Eqs. (B1) and (B2). What remains is to compute the time-dependent matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$, and $\rho^{(d)}_{k,22}(t)$ that appear in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) (with the suppressed time argument reinstated). For the noiseless pure ensemble, this can be done by numerically integrating the von Neumann equation in the diagonal basis,
$\dot{\rho}_{0,k}(t)=-i[H_{0,k}(t),\rho_{0,k}(t)],$ (B41)
for all $k$, with $t$ the time when the quench is completed. As initial condition we impose that the system is in the noiseless ground state at time $t_{i}$, i.e., that all modes $k$ occupy the lower levels $|\phi_{k}^{-}(t_{i})\rangle$. As expected for a pure ensemble, note that the matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$ and $\rho^{(d)}_{k,22}(t)$ defined in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) (with the time argument suppressed) map onto those in Eq. (III), where in the latter equation the density operator for mode $k$ is denoted $\rho_{0,k}(t)$ while in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) the more general symbol $\rho_{k}$ - adaptable to pure as well as mixed ensembles - is used.
@@
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}[\sigma^{z},[\sigma^{z},\rho_{k}(t)]],$ (B42)
where $\rho_{k}(t)=\langle\rho_{\eta,k}(t)\rangle$ is the averaged density matrix for mode $k$ in the presence of white noise, with the average taken over the white-noise distribution $\{\eta(t)\}$ specified in Eq. (10). By numerically integrating the master equation, Eq. (B42), for all $k$, with the same initial condition as in the noiseless case, we extract the matrix elements $\rho^{(d)}_{k,11}(t),\rho^{(d)}_{k,12}(t),\rho^{(d)}_{k,21}(t)$ and $\rho^{(d)}_{k,22}(t)$ in the diagonal basis. In this way we finally obtain the mean concurrences $\langle C_{l,l+1}(t)\rangle$ and $\langle C_{l,l+2}(t)\rangle$ for nearest- and next-nearest-neighbor spins.
@@
Appendix C: Noise-averaged nonequilibrium concurrence: exact master equation
@@
Let $\rho(t)=\langle\rho_{\eta}(t)\rangle$ be the ensemble average over many noise realizations (all with a common noise-free initial condition). The averaged von Neumann equation (C2) then takes the form
$\dot{\rho}(t)=-i[H_{0}(t),\rho(t)]-i[H_{1}(t),\langle\eta(t)\rho_{\eta}(t)\rangle].$ (C4)
@@
$\langle\eta(t)\rho_{\eta}(t)\rangle=\langle\eta(t)\rangle\langle\rho_{\eta}(t)\rangle+\int_{0}^{t}\langle\eta(t)\eta(s)\rangle\langle\frac{\delta\rho_{\eta}}{\delta\eta}\rangle ds,$ (C5)
@@
$\frac{\delta\rho_{\eta}}{\delta\eta}=\frac{\partial\dot{\rho}_{\eta}}{\partial\eta}-\frac{\mbox{d}}{\mbox{d}t}\frac{\partial\dot{\rho}_{\eta}}{\partial\dot{\eta}}.$ (C6)
Combining Eqs. (C6), (C1), and (C2) gives
$\frac{\delta\rho_{\eta}}{\delta\eta}=-i[H_{1}(t),\eta(t)].$ (C7)
The master equation follows by inserting Eq. (C5) with $\langle\eta(t)\rangle=0$ into (C4), using Eqs. (10) and (C7),
$\dot{\rho}(t)=-i[H_{0}(t),\rho(t)]-\frac{\xi^{2}}{2}\Big[H_{1}(t),[H_{1}(t),\rho(t)]\Big],$ (C8)
where the factor of $1/2$ multiplying the noise term comes from the delta function in Eq. (10) having support at the edge of the integral in Eq. (C5).
As follows from the discussion in Sec. II, the TFI Hamiltonian $H_{0}(t)$ in Eq. (1) with an added white-noise term gets expressed as a sum over decoupled mode Hamiltonians $H_{k}(t)=H_{0,k}(t)+\eta(t)H_{1}$ after a Jordan-Wigner transformation, with $H_{0,k}(t)$ given in Eq. (4) and with $H_{1}=\sigma^{z}$. As a consequence, the density matrix $\rho_{\eta}(t)$ corresponding to $H_{0}(t)$ with the added white-noise term has a direct product structure, i.e., $\rho_{\eta}(t)=\otimes_{k}\rho_{k,\eta}(t)$ with the $2\times 2$ density matrix $\rho_{k,\eta}(t)$ satisfying $\dot{\rho}_{k,\eta}(t)=-i[H_{0,k}(t),\rho_{k,\eta}(t)]$ for a single realization of the noise function $\eta$, common to all modes $k$. The master equation for the noise-averaged density matrix $\rho_{k}(t)=\langle\rho_{k,\eta}(t)\rangle$ is read off from Eq. (C8), using that $H_{1}=\sigma^{z}$, with the result
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}\Big[\sigma^{z},[\sigma^{z},\rho_{k}(t)]\Big].$ (C9)